a stone is projected vertically upwards from the ground to a height of 25m.if it is reaches the point of projection ofter some time,then the distance and displacements of the stone are

Grade:12th pass

2 Answers

noogler
8 years ago
distance=25+25=50m
displacement=25+(-25)=0m
 
displacement is a vector so we should consider its direction. let upward motion of stone is take +ve then downward motion of stone will be -ve
distance is a scalar so we donot consider its direction
erra akhil
8 years ago
Dear Chandrasekhar,
Distance literally means how much long the obect has travelled.
Displacent literally means how far out of a place an object is with respect to the original position.
Here,displacent=0,because the object position is not changed.
distance=25+25=50; because the bbody travels 25m to go up and then 25 to come down.
Approve if my answer helped you.